A dam is put on a river that has a unblocked flow rate of 2 meter/second. The new dam then blocks the water and forces it through a generation system. If 1,000,000 cubic meters of water flowed each hour in the river before the dam went in, how much electric power ( Energy / time) can be expected to be grabbed from this water once the dam is in place? Use a 50% effective conversion of the kinetic energy of the water flow to electricity.
Answer #1

The kinetic energy is given by K.E. = (1/2)mv2

K.E. = (1/2)*(1000000*1000Kg/hour)*(2m/s)2

K.E. = 2 x 109 J/hour

As electricity produced is 50% of kinetic energy, Electric Energy = 0.5*K.E.

Hence, Electric Energy produced = 109 J/hour

I hope i was able to answer your question. If you need anymore explanation, please ask in comments. I would be happy to help.

Please rate the solution if you are satisfied with the solution.